Green light at 546 nm is diffracted by a grating with 3000 lines per cm. The light is normally incident on the diffraction grating.Through what angle is the light diffracted in the first order? Through what angle is the light diffracted in the fifth order?
Answer #1

Given,

Wavelength = 546nm
Slit Spacing (d) = 1/3000 cm = 3.33 * 10^-6


We know -
d*sin(theta) = m*wavelength

For Part a)
Angle, light diffracted in the first order =

3.33 * 10^-6 * sin(theta) = 1 * 546 * 10^ -9
theta = 9.4360

Angle, light diffracted in the first order = 9.4360

B) Angle, light diffracted in the fifth order =
3.33 * 10^-6 * sin(theta) = 5 * 546 * 10^ -9

Angle, light diffracted in the fifth order = 55.070